p=14,000, m = 14, r = 6% compounded annually
Answer by mananth(16949)

Principal P = 14000
Amount= A
years=n 14
compounded 1 times a year
Rate = 6.00 0.06
Amount =P*((n+r)/n)^n

Amount =14000*(1+0.06)^14*1
Amount =14000*(1+0.06)^14
14000*(1.06)^14
Amount =31652.66